# Classification of compact manifolds with positive isotropic curvature

### v6

@@ 2 $\varepsilon$-caps and ancient $\kappa$-solutions
(cf. [33], Corollaries 6.20 and 6.22 in [7], Theorem 3.10 in [10], Proposition 2.9 in [19])   Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$, $C_{2}=C_{2}(n,\theta,\varepsilon)$ and $C_{3}=C_{3}(n,\theta)$ with the following property: Suppose $(\mathcal{O},g(t))$ is an orbifold ancient $\kappa$-solution of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$. Then either $(\mathcal{O},g(t))$ is compact and strictly PIC2 for any $t$ (hence diffeomorphic to a spherical orbifold), or for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ of $\mathcal{O}$ with $\overline{B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t_{0}$ lying (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$, which falls into one of the following two categories:
(a) $U$ is a strong $\varepsilon$-neck centered at $(x_{0},t_{0})$,
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$;
@@
Proof.    One can argue as in the proof of Proposition 2.9 in [19], using Propositions 2.4 and 2.6. $\Box$
@@
(a) $U$ is an $\varepsilon$-neck centered at $(x_{0},t_{0})$, and on $U\times[t_{0}-\min\{R(x_{0},t_{0})^{-1},t_{0}\},t_{0}]$ the solution $\hat{g}_{\Gamma}(t)$ is, after scaling with $R(x_{0},t_{0})$ and shifting $t_{0}$ to zero, $\varepsilon$-close to the corresponding subset of the evolving round cylinder $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ over the time interval $[-\min\{t_{0}R(x_{0},t_{0}),1\},0]$ with scalar curvature 1 at the time zero, and $U$ is disjoint from the surgery cap (in particular, $x_{0}\notin B(p_{\Gamma},0,\varepsilon^{-1})$) when $t_{0}\leq R(x_{0},t_{0})^{-1}$, or
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$ (diffeomorphic to $\mathbb{R}^{n}//\Gamma$);
[…]
@@ 3 Existence of $(r,\delta)$-surgical solutions and applications
(cf. Proposition 3.2 in [19])   Let $0<\varepsilon_{0}\leq\varepsilon_{3}$, where the constant $\varepsilon_{3}$ is as in Proposition 5.7 in [19]. Fix $0<\varepsilon<\tilde{\varepsilon}_{1}(\varepsilon_{0})$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma 5.6 in [19]. Let $(\mathcal{O},g)$ be a closed, connected $n$-orbifold with at most isolated singularities. Suppose that each point of $\mathcal{O}$ is a center of an $\varepsilon$-neck or an $\varepsilon$-cap. Then $\mathcal{O}$ is diffeomorphic to a spherical orbifold or a connected sum of at most two spherical orbifolds with at most isolated singularities. If we assume in addition that $\mathcal{O}$ is a manifold, then either $\mathcal{O}$ is diffeomorhic to a spherical space form, or $\mathcal{O}$ is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries.
@@
If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $\mathbb{R}^{n}//\Gamma$ ($|\Gamma|\geq 2$) but no $\varepsilon$-caps of the other types, by using the definition of $\varepsilon$-caps, Lemma 5.6 and Proposition 5.7 in [19], and Theorem 1.5 in Chapter 8 of [17], we see that it is diffeomorphic to $D^{n}//\Gamma\cup_{h}D^{n}//\Gamma$, where $h\in\text{Isom}(\mathbb{S}^{n-1}/\Gamma)$, so it is diffeomorphic to a connected sum $\mathbb{S}^{n}//\Gamma\sharp\mathbb{S}^{n}//\Gamma$ (where the connected sum occurs at two singular points). In fact, it is diffeomorphic to $\mathbb{S}^{n}//\Gamma$; cf. Theorem 2.2 in Chapter 8 of [17] and Proposition 7.7 in [2].
@@
(2) $f_{0}$ is smooth, concave and positive on $(-n,n)$;
@@
Following Chapter 12 of [27], we argue the existence of the function $f_{0}$. Let $h_{0}(s)=\frac{s-1}{\sqrt{(n-1)^{2}-(s-1)^{2}}}$ for $s\in[0,n)$, and $\lambda:(-\infty,\infty)\rightarrow[0,1]$ be a non-decreasing smooth function which is identically 0 on $(-\infty,n-\frac{1}{2}+\nu]$ and identically 1 on $[n-\nu,\infty)$ for some positive number $\nu<\frac{1}{10}$. Note that
[…]
@@
[…]
@@
[…]
be the revolutionary hypersurface in $\mathbb{R}^{n+1}$ obtained via rotating the graph of the function $f_{0}$; compare Chapter 12 in [27]. Then $\Sigma(f_{0})$ has an obvious $O(n)\times\mathbb{Z}_{2}$ symmetry, and $\Sigma(f_{0})$ is $O(n)\times\mathbb{Z}_{2}$-equivariantly diffeomorphic to the standard sphere $\mathbb{S}^{n}$. It follows that the spherical orbifold $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$ is diffeomorphic to the union of certain Riemannian manifold-with-boundary isometric to some $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}[-1,1]$ (which is the quotient of $\mathbb{S}^{n-1}/\Gamma\times[-1,1]$ by the ${\mathbb{Z}_{2}}$-action generated by $\hat{\sigma}$ with $\hat{\sigma}(x,s)=(\sigma(x),-s)$ for $x\in\mathbb{S}^{n-1}/\Gamma$ and $s\in[-1,1]$) and a “standard” orbifold cap (which is the quotient of a “standard” (rotationally symmetric) cap (which has an end isometric to $\mathbb{S}^{n-1}\times[0,\mu]$ for some small $\mu>0$) by $\Gamma$) along their boundaries via an isometry between the boundaries. By combining this observation with the argument on p. 105 of [8] and Proposition 2.1, we see that any connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{1}\rangle\sharp\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{2}\rangle$ (here $\Gamma$ may be trivial; when $\Gamma$ is nontrivial we require that the connected sum occurs at two singular points) is diffeomorphic to the union of two Riemannian manifolds-with-boundary along their boundaries via an isometry between the boundaries, where each of the Riemannian manifold-with-boundary is isometric to some $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}[-1,1]$. (When $\Gamma$ is trivial, we use the Cerf-Palais disk theorem to replace Proposition 2.1.) Such a union has $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry. (By the way, similarly, by using Proposition 2.1 and (the proof of) Theorem 2.2 in Chapter 8 of [17] (cf. also Proposition 7.7 in [2] and Theorem 10.2 in [20]) one can show that any connected sum $\mathbb{S}^{n}//\Gamma\sharp\mathbb{S}^{n}//\Gamma$ (where the connected sum occurs at two singular points when $|\Gamma|\geq 2$) is diffeomorphic to $\mathbb{S}^{n}//\Gamma$, and any connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\sharp\mathbb{S}^{n}//\Gamma$ (where the connected sum occurs at two singular points when $|\Gamma|\geq 2$) is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$.)
@@
We fix $\varepsilon_{0}=\min\{\varepsilon_{3},\varepsilon_{4}\}$, where the constant $\varepsilon_{3}$ is as in Proposition 5.7 in [19], and the constant $\varepsilon_{4}$ is as in Proposition 2.4. Choose $0<\varepsilon<\frac{1}{4}\min\{\frac{1}{10^{3}n},\tilde{\varepsilon}_{1}(\varepsilon_{0})\}$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma 5.6 in [19].
@@
$\mathcal{F}_{t}\subset\{\text{Rm}\hskip 2.84526pt|\hskip 2.84526pt\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}\}\cap\{\text{Rm}\hskip 2.84526pt|\hskip 2.84526pt\text{Rm}+f(R)\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}2}\}$
for any $t\in[0,\hat{T}]$, where $f:[0,\infty)\rightarrow[0,\infty)$ is an increasing concave function satisfying $\text{lim}_{s\rightarrow\infty}\frac{f(s)}{s}=0$, and $\theta$ is a positive number. As in [7], note that here $f$ and $\theta$ depend only on the initial data $(\mathcal{O},g_{0})$. By (a version of) Hamilton’s maximum principle the curvature tensor of $(\mathcal{O},g(t))$ lies in $\mathcal{F}_{t}$ for any $t\in[0,T)$. Let $\beta=\beta(\varepsilon)$ be the constant given in Lemma 5.8 in [19], and choose
@@
where the constants on the RHS are from Propositions 2.7 and 2.8.
@@
As in [19], with the help of Brendle’s curvature pinching estimates (see Theorem 1.2 and Corollary 1.3 in [7]), and Propositions 2.7, 2.8, and 3.1, we can show the following theorem.
@@
(cf. Theorem 11.2 in [7], Theorem 4.6 in [10], and Theorem 4.11 in [19])    Let $(\mathcal{O},g_{0})$ be a compact Riemannian orbifold of dimension $n\geq 12$ with at most isolated singularities and with positive isotropic curvature. Let $\varepsilon$ and $C$ be chosen as above. Then we can find positive numbers $\hat{r}$ and $\hat{\delta}$ such that there exists an $(\hat{r},\hat{\delta})$-surgical solution (with $(4\varepsilon,4C)$-control) to the Ricci flow starting with $(\mathcal{O},g_{0})$, which becomes extinct in finite time.
@@
First we undo the connected sum in the orbifold components which are diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\hskip 2.84526pt\sharp\hskip 2.84526pt\mathbb{S}^{n}//(x,\pm x^{\prime})$ (where the connected sum occurs at two regular points) or $\mathbb{S}^{n}//(x,\pm x^{\prime})\hskip 2.84526pt\sharp\hskip 2.84526pt\mathbb{R}P^{n}$, and do orbifold connected sum operations among the orbifolds thus gotten and all other orbifold components to resolve the orbifold singularities of the components, which are created by the Ricci flow surgeries. (If a component is a smooth manifold, we don’t need to worry about it at this step.) By using the proof of Proposition 3.1 we see that what we get in this way are some manifolds, denoted by $Y_{j}$, where each of them is diffeomorphic to either a connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{1}\rangle\sharp\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{2}\rangle$ or a self-connected sum of $\mathbb{S}^{n}//\Gamma$ at the two singular points. (Note that $\mathbb{S}^{n}//(x,\pm x^{\prime})$ is isometric to certain $\mathbb{S}^{n}//\Gamma$, where $\Gamma\cong\mathbb{Z}_{2}<O(n)$ acts on $\mathbb{S}^{n-1}$ antipodally.) By the last claim in the proof of Proposition 3.1 in the former case the resulting manifold admits $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry. By an argument similar to that in the last claim in the proof of Proposition 3.1, we see that any self-connected sum of $\mathbb{S}^{n}//\Gamma$ (where $|\Gamma|\geq 2$) at the two singular points must also admit $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry. (By the way, a similar argument with Proposition 2.1 replaced by Theorem 3.2 in Chapter 8 of [17] (see also Corollary 3.7 in Chapter III of [22] and Proposition 2.1 in [18]) shows that any self-connected sum of $\mathbb{S}^{n}$ also admits $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry; but this is well-known.)
@@
Proof of Corollary 1.2. Let $M$ satisfy the assumptions of Corollary 1.2. Then $M$ has a connected sum decomposition as in the conclusion of Theorem 1.1. Note that $n$ is even and $M$ is orientable by assumption, the only orientable $n$-dimensional spherical space form is the sphere $\mathbb{S}^{n}$, and a manifold-with-boundary of the form $\mathbb{S}^{n-1}/\Gamma_{i}\times_{\mathbb{Z}_{2}}[-1,1]$ is non-orientable (recall that for even $n$ the action of any finite subgroup of $O(n)$ on $\mathbb{S}^{n-1}$ must be orientation-preserving if it is free), hence a manifold which is diffeomorphic to the union of two manifolds-with-boundary both of the form $\mathbb{S}^{n-1}/\Gamma_{i}\times_{\mathbb{Z}_{2}}[-1,1]$ along the boundaries is also non-orientable, so we can assume that in the connected sum decomposition of $M$ each summand is diffeomorphic to the total space of an orientable $\mathbb{S}^{n-1}/\Gamma_{j}$-bundle over $\mathbb{S}^{1}$ with structure group $\text{Isom}(\mathbb{S}^{n-1}/\Gamma_{j})$. By comparing the first Betti numbers we see that there is only one such summand, and we are done. $\Box$

### v8

@@ 2 $\varepsilon$-caps and ancient $\kappa$-solutions
(cf. [43], Corollaries 6.20 and 6.22 in [11], Theorem 3.10 in [15], Proposition 2.9 in [27]) Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$, $C_{2}=C_{2}(n,\theta,\varepsilon)$ and $C_{3}=C_{3}(n,\theta)$ with the following property: Suppose $(\mathcal{O},g(t))$ is an orbifold ancient $\kappa$-solution of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}\text{id}\in C_{\text{PIC}}$. Then either $(\mathcal{O},g(t))$ is compact and strictly PIC2 for any $t$ (hence diffeomorphic to a spherical orbifold), or for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ of $\mathcal{O}$ with $\overline{B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$ and with the values of the scalar curvature in $U$ at time $t_{0}$ lying (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$, which falls into one of the following two categories:
(a) $U$ is a strong $\varepsilon$-neck diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ centered at $(x_{0},t_{0})$, or
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$ with any cross-section of the $\varepsilon$-neck contained in the end of it diffeomorphic to $\mathbb{S}^{n-1}/\Gamma$
for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$, and
$vol_{g(t_{0})}(U)>(C_{2}|\Gamma|)^{-1}R(x_{0},t_{0})^{-n/2};$
@@
Proof. One can argue as in the proof of Proposition 2.9 in [27], using Propositions 2.5 and 2.7. $\Box$
@@
(a) $U$ is an $\varepsilon$-neck diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ centered at $(x_{0},t_{0})$, and on $U\times[t_{0}-\min\{R(x_{0},t_{0})^{-1},t_{0}\},t_{0}]$ the solution $\hat{g}_{\Gamma}(t)$ is, after scaling with $R(x_{0},t_{0})$ and shifting $t_{0}$ to zero, $\varepsilon$-close to the corresponding subset of the evolving round cylinder $\mathbb{S}^{n-1}/\Gamma\times\mathbb{R}$ over the time interval $\min\{t_{0}R(x_{0},t_{0}),1\},0]$ with scalar curvature 1 at the time zero, and $U$ is disjoint from the surgery cap (in particular, $x_{0}\notin B(p_{\Gamma},0,\varepsilon^{-1})$, where $p_{\Gamma}$ denotes the tip of the orbifold standard solution) when $t_{0}\leq R(x_{0},t_{0})^{-1}$, or
(b) $U$ is an $\varepsilon$-cap diffeomorphic to $\mathbb{R}^{n}//\Gamma$ centered at $(x_{0},t_{0})$,
and $vol_{\hat{g}_{\Gamma}(t_{0})}(U)>(C_{2}^{\prime}|\Gamma|)^{-1}R(x_{0},t_{0})^{-n/2}$; moreover, the scalar curvature in $U$ at time $t_{0}$ satisfies the derivative estimates
$|\nabla R|<C_{3}^{\prime}R^{\frac{3}{2}}\hskip 22.76219pt\mbox{and}\hskip 22.76219pt|\frac{\partial R}{\partial t}|<C_{3}^{\prime}R^{2}.$
@@ 3 Existence of $(r,\delta)$-surgical solutions and applications
(cf. Proposition 3.2 in [27]) Let $0<\varepsilon_{0}\leq\varepsilon_{3}$, where the constant $\varepsilon_{3}$ is as in Proposition A.7 in [27]. Fix $0<\varepsilon<\tilde{\varepsilon}_{1}(\varepsilon_{0})$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma A.6 in [27]. Let $(\mathcal{O},g)$ be a closed, connected $n$-orbifold with at most isolated singularities. Suppose that each point of $\mathcal{O}$ is a center of an $\varepsilon$-neck or an $\varepsilon$-cap. Then $\mathcal{O}$ is diffeomorphic to a spherical orbifold or a connected sum of at most two spherical orbifolds with at most isolated singularities. If we assume in addition that $\mathcal{O}$ is a manifold, then either $\mathcal{O}$ is diffeomorhic to a spherical space form, or $\mathcal{O}$ is diffeomorphic to a quotient manifold of $\mathbb{S}^{n-1}\times\mathbb{R}$ by standard isometries.
@@
If $\mathcal{O}$ contains (at least) an $\varepsilon$-cap of type $\mathbb{R}^{n}//\Gamma$ ($|\Gamma|\geq 2$) but no $\varepsilon$-caps of the other types, by using the definition of $\varepsilon$-caps, Lemma A.6 and Proposition A.7 in [27], and Theorem 1.5 in Chapter 8 of [25], we see that it is diffeomorphic to the double of $D^{n}//\Gamma$, hence it is diffeomorphic to $\mathbb{S}^{n}//\Gamma$.
@@
(2) $f_{0}$ is smooth, concave and positive on (n,n)(-n,n);
@@
Following Chapter 12 of [37], we argue the existence of the function $f_{0}$. Let $h_{0}(s)=\frac{s-1}{\sqrt{(n-1)^{2}-(s-1)^{2}}}$ for $s\in[0,n)$, and $\lambda:(-\infty,\infty)\rightarrow[0,1]$ be a non-decreasing smooth function which is identically 0 on $(-\infty,n-\frac{1}{2}+\nu]$ and identically 1 on $n-\nu,\infty)$ for some positive number $\nu<\frac{1}{10}$. Note that
$\int_{1}^{n}h_{0}(s)\lambda(s+n-\frac{3}{2})ds>\int_{3/2}^{n}h_{0}(s)ds>\sqrt{(n-1)(n-2)},$
@@
$f_{0}(s)=\int_{s}^{n}h_{0}(r)\lambda(r+s_{0})dr\hskip 2.84526pt\text{for}\hskip 2.84526pts\in[0,n]\hskip 5.69054pt\text{and}\hskip 5.69054ptf_{0}(s)=f_{0}(-s)\hskip 2.84526pt\text{for}\hskip 2.84526pts\in[-n,0],$
@@
$\Sigma(f_{0})=\{(x_{1},x_{2},\cdot\cdot\cdot,x_{n+1})\in\mathbb{R}^{n+1}\hskip 2.84526pt|\hskip 2.84526pt-n\leq x_{n+1}\leq n,\hskip 5.69054ptx_{1}^{2}+x_{2}^{2}+\cdot\cdot\cdot+x_{n}^{2}=f_{0}(x_{n+1})^{2}\}$
be the revolutionary hypersurface in $\mathbb{R}^{n+1}$ obtained via rotating the graph of the function $f_{0}$; compare Chapter 12 in [37]. Then $\Sigma(f_{0})$ has an obvious $O(n)\times\mathbb{Z}_{2}$ symmetry, and $\Sigma(f_{0})$ is $O(n)\times\mathbb{Z}_{2}$-equivariantly diffeomorphic to the standard sphere $\mathbb{S}^{n}$. It follows that the spherical orbifold $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$ is diffeomorphic to the union of a certain Riemannian manifold-with-boundary isometric to some $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}[-1,1]$ (which is the quotient of $\mathbb{S}^{n-1}/\Gamma\times[-1,1]$ by the $\mathbb{Z}_{2}$}-action generated by $\hat{\sigma}$ with $\hat{\sigma}(x,s)=(\sigma(x),-s)$ for $x\in\mathbb{S}^{n-1}/\Gamma$ and $s\in[-1,1]$) and a “standard” orbifold cap (which is the quotient of a “standard” (rotationally symmetric) cap (which has an end isometric to $\mathbb{S}^{n-1}\times[0,\mu]$ for some small $\mu>0$) by $\Gamma$) along their boundaries via an isometry between the boundaries. By combining this observation with the argument on p. 105 of [12] and Proposition 2.2, we see that any connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{1}\rangle\sharp\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{2}\rangle$ (here $\Gamma$ may be trivial; when $\Gamma$ is nontrivial we require that the connected sum occurs at two singular points) is diffeomorphic to the union of two Riemannian manifolds-with-boundary along their boundaries via an isometry between the boundaries, where each of the Riemannian manifold-with-boundary is isometric to some $\mathbb{S}^{n-1}/\Gamma\times_{\mathbb{Z}_{2}}[-1,1]$. (When $\Gamma$ is trivial, we use the Cerf-Palais disk theorem to replace Proposition 2.2.) Such a union has $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry. (By the way, similarly, by using Proposition 2.2 and (the proof of) Theorem 2.2 in Chapter 8 of [25] (cf. also Proposition 7.7 in [3] and Theorem 10.2 in [28]) one can show that any connected sum $\mathbb{S}^{n}//\Gamma\sharp\mathbb{S}^{n}//\Gamma$ (where the connected sum occurs at two singular points when $|\Gamma|\geq 2$) is diffeomorphic to $\mathbb{S}^{n}//\Gamma$, and any connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\sharp\mathbb{S}^{n}//\Gamma$ (where the connected sum occurs at two singular points when $|\Gamma|\geq 2$) is diffeomorphic to $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle$.)
@@
We fix $\varepsilon_{0}=\min\{\varepsilon_{3},\varepsilon_{4}\}$, where the constant $\varepsilon_{3}$ is as in Proposition A.7 in [27], and the constant $\varepsilon_{4}$ is as in Proposition 2.5. Choose $0<\varepsilon<\frac{1}{4}\min\{\frac{1}{10^{4}n},\tilde{\varepsilon}_{1}(\varepsilon_{0})\}$, where $\tilde{\varepsilon}_{1}(\cdot)$ is as in Lemma A.6 in [27].
@@
$\mathcal{F}_{t}\subset\{\text{Rm}\hskip 2.84526pt|\hskip 2.84526pt\text{Rm}-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}\text{id}\in C_{\text{PIC}}\}\cap\{\text{Rm}\hskip 2.84526pt|\hskip 2.84526pt\text{Rm}+f(R)\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\vtop{\halign{#\cr$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}}{}}{\mathbin{\vtop{\halign{#\cr$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}}{}}\text{id}\in C_{\text{PIC}2}\}$
for any $t\in[0,\hat{T}$], where $f:[0,\infty)\rightarrow[0,\infty)$ is an increasing concave function satisfying $\text{lim}_{s\rightarrow\infty}\frac{f(s)}{s}=0$, and $\theta$ is a positive number. As in [11], note that here $f$ and $\theta$ depend only on the initial data $(\mathcal{O},g_{0})$. By (a version of) Hamilton’s maximum principle the curvature tensor of $(\mathcal{O},g(t))$ lies in $\mathcal{F}_{t}$ for any $t\in[0,T)$. Let $\beta=\beta(\varepsilon)$ be the constant given in Lemma A.8 in [27], and choose
@@
where the constants on the RHS are from Propositions 2.8 and 2.9.
@@
As in [27], with the help of Brendle’s curvature pinching estimates (see Theorem 1.2 and Corollary 1.3 in [11]), and Propositions 2.8, 2.9, and 3.1, we can show the following theorem.
@@
[…]
@@
First we undo the connected sum in the orbifold components which are diffeomorphic to $\mathbb{S}^{n}//(x,\pm x^{\prime})\hskip 2.84526pt\sharp\hskip 2.84526pt\mathbb{S}^{n}//(x,\pm x^{\prime})$ (where the connected sum occurs at two regular points) or $\mathbb{S}^{n}//(x,\pm x^{\prime})\hskip 2.84526pt\sharp\hskip 2.84526pt\mathbb{R}P^{n}$, and do orbifold connected sum operations among the orbifolds thus gotten and all other orbifold components to resolve the orbifold singularities of the components, which are created by the Ricci flow surgeries. (If a component is a smooth manifold, we don’t need to worry about it at this step.) By using the proof of Proposition 3.1 we see that what we get in this way are some manifolds, denoted by $Y_{j}$, where each of them is diffeomorphic to either a connected sum $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{1}\rangle\sharp\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}_{2}\rangle$ or a self-connected sum of $\mathbb{S}^{n}//\Gamma$ at the two singular points. (Note that $\mathbb{S}^{n}//(x,\pm x^{\prime})$ is isometric to a certain $\mathbb{S}^{n}//\Gamma$, where $\Gamma\cong\mathbb{Z}_{2}<O(n)$ acts on $\mathbb{S}^{n-1}$ antipodally and then on $\mathbb{S}^{n}$ by suspension.) By the last claim in the proof of Proposition 3.1 in the former case the resulting manifold admits $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry. By an argument similar to that in the last claim in the proof of Proposition 3.1, we see that any self-connected sum of $\mathbb{S}^{n}//\Gamma$ (where $|\Gamma|\geq 2$) at the two singular points must also admit $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry. (By the way, a similar argument with Proposition 2.2 replaced by Theorem 3.2 in Chapter 8 of [25] (see also Corollary 3.7 in Chapter III of [30] and Proposition 2.1 in [26]) shows that any self-connected sum of $\mathbb{S}^{n}$ also admits $\mathbb{S}^{n-1}\times\mathbb{R}$-geometry; but this is well-known.)
@@
Proof of Corollary 1.2. Let $M$ satisfy the assumptions of Corollary 1.2. Then $M$ has a connected sum decomposition as in the conclusion of Theorem 1.1. Note that $n$ is even and $M$ is orientable by assumption, the only orientable $n$-dimensional spherical space form is the sphere $\mathbb{S}^{n}$, and a manifold-with-boundary of the form $\mathbb{S}^{n-1}/\Gamma_{i}\times_{\mathbb{Z}_{2}}[-1,1]$ is non-orientable (recall that for even $n$ the action of any finite subgroup of $O(n)$ on $\mathbb{S}^{n-1}$ must be orientation-preserving if it is free), hence a manifold which is diffeomorphic to the union of two manifolds-with-boundary both of the form $\mathbb{S}^{n-1}/\Gamma_{i}\times_{\mathbb{Z}_{2}}[-1,1]$ along the boundaries is also non-orientable, so we can assume that in the connected sum decomposition of $M$ each summand is diffeomorphic to the total space of an orientable $\mathbb{S}^{n-1}/\Gamma_{j}$-bundle over $\mathbb{S}^{1}$ with structure group $\text{Isom}(\mathbb{S}^{n-1}/\Gamma_{j})$. By comparing the first Betti number we see that there is only one such summand, and we are done. $\Box$